OpenStudy (anonymous):

Find the area of a circle with a diameter of 16 inches. Use 3.14 for pi. (1 point) Select one: a. 50.24 in2 b. 100.48 in2 c. 200.96 in2 d. 251.2 in2

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

Area of circle = pi * (diameter)^2/4

OpenStudy (anonymous):

i w it would have a formula for it

OpenStudy (anonymous):

i got C all i really needed was the formula thank you so much:)

OpenStudy (anonymous):

The volume of a fish tank is 40 cubic feet. If the density is 0.4 fish/feet^3, how many fish are in the tank? (1 point) Select one: a. 16 b. 4 c. 100 d. 10

OpenStudy (anonymous):

there has to be a formula for this one to

OpenStudy (anonymous):

@HB0102

hartnn (hartnn):

density = number / volume

hartnn (hartnn):

0.4= x/40

OpenStudy (anonymous):

so what am i supposed to do

hartnn (hartnn):

find x multiply 40 on both sides

OpenStudy (anonymous):

0.4= x/40 how are we supposed to multiply 40 on the right side

hartnn (hartnn):

left side 0.4*40 right side x*40/40 = x

OpenStudy (anonymous):

what does x equal

hartnn (hartnn):

0.4*40

hartnn (hartnn):

x is the number of fish in the tank

OpenStudy (anonymous):

but you said right side x*40/40 = x

hartnn (hartnn):

thts correct. so?

OpenStudy (anonymous):

but i thought left side 0.4*40

hartnn (hartnn):

it is.

hartnn (hartnn):

why and where are you confused ?

hartnn (hartnn):

calculate whats 0.4*40 thats would be your answer

OpenStudy (anonymous):

The density of a fish tank is 0.4 fish/feet^3 There are 12 fish in the tank. What is the volume of the tank? (1 point) Select one: a. 3 ft3 b. 30 ft3 c. 48 ft3 d. 96 ft3

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

same formula

hartnn (hartnn):

density = number / volume 0.4 = 12/ v

OpenStudy (anonymous):

i know i did 12 times 0.4 and 4.8 came out

hartnn (hartnn):

no.. 0.4 = 12/v so v = 12/0.4 = 120/4

OpenStudy (anonymous):

so do 12/4 then

OpenStudy (anonymous):

i mean 120/4

OpenStudy (anonymous):

i got 30 is that correct

hartnn (hartnn):

yes

OpenStudy (anonymous):

thank you